Define a partial differential equation and give an example.

Hence find the differential equation arising from F[y/x,z/x3]=0


1
Expert's answer
2021-11-03T18:18:17-0400

The partial differential equation is an equation which imposes relations between the various partial derivatives of a multivariable function. For exaple, "\frac{d^2u}{d x^2}-\frac{d^2u}{d y^2}=0."

Let's find the differential equation arising from "F[y\/x,z\/x3]=0" .

Let "f(x;y;z)=F\left(\frac{y}{x};\frac{z}{x^3}\right)."

"f'(x)=-\frac{y}{x^2}F'_{x_1}-3\frac{z}{x^4}F'_{x_1}; f'(y)=F'_{x_1}/x; f'(z)=F'_{x_1}/x^3." So we have "F'_{x_1}=xf'(x); F'_{x_2}=x^3f'(z)" . Since the function "x(y;z)" can be given by from the equality "f(x;y;z)=0" by the implicit function theorem. We have "x'(y)=-\frac{f'(y)}{f'(x)}, x'(z)=-\frac{f'(z)}{f'(x)}." From the equality "f'(x)=-\frac{y}{x^2}F'_{x_1}-3\frac{z}{x^4}F'_{x_1}" we have "1=\frac{y}{x}x'(y)+3\frac{z}{x}x'(z)" or "yx'(y)+3zx'(z)=x."